OpenStudy (cyclone3260):

Multiply and Simplify ³√5a2 * ³√2a Step by step please?

jhonyy9 (jhonyy9):

do you know how you can multiple radicals ?

jhonyy9 (jhonyy9):

so there are cuberoot(5a^2) *cuberoot(2a) = ? yes ?

jhonyy9 (jhonyy9):

for example than you know that sqrt(a) *sqrt(b) = sqrt(a*b) do you can making now you above wrote exercise ?

jhonyy9 (jhonyy9):

sqrt mean squarroot just in case of your exercise there are cuberoot

jhonyy9 (jhonyy9):

ok i will give you one example using cuberoot cuberoot(a) *cuberoot(b) = cuberoot(a*b) hope this helped

jhonyy9 (jhonyy9):

\[\sqrt[3]{5a^2} *\sqrt[3]{2a} = \sqrt[3]{5a^2 *2a} = ?\]